OpenStudy (anonymous):

Find the total resistance in a parallel dc circuit with two branches rated at 4 and 18, respectively.

OpenStudy (anonymous):

resistances are in parallel. \[\frac{ 1 }{ R }=\frac{ 1 }{ R _{1} }+\frac{ 1 }{ R _{2} }\]